Find the point-slope equations of the line using the point (3, 5) and slope 3/2 .

Answer by Alan3354(69443)

Do it like this one.
You already know the slope, so start there.
Email me via the thank you not to check your work, or for assistance.
------------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.